OpenStudy (some.random.cool.kid):

The complex solution to a quadratic equation is x equals start fraction negative 10 plus or minus square root of negative 700 end square root over two end fraction full stop Write this solution in standard form, x = a ± bi, where a and b are real numbers. Justify your answer by showing your work.

OpenStudy (some.random.cool.kid):

need help bad

OpenStudy (some.random.cool.kid):

sqrt of the -700 is 26.4 and now im stuck

OpenStudy (some.random.cool.kid):

thats the image I just attached the file instead because the writing was to messy.

Nnesha (nnesha):

don't use calculator forsqrt{-700} there is a negative sign so you should get imaginary solution so factor out the -700 what is the largest factor of 700 that should be the perfect square root

OpenStudy (some.random.cool.kid):

350?

OpenStudy (some.random.cool.kid):

idk a perfect sqrt for 700 close ive got is 26.4 and thats a decimal

Nnesha (nnesha):

hmm 350 isn't a perfect square root what two numbers would you multiply to get 700 ??

Nnesha (nnesha):

here are some example of perfect square root sqrt{16}= 4 sqrt{25}=5

OpenStudy (some.random.cool.kid):

350and 2?

OpenStudy (some.random.cool.kid):

that can get you 700?

OpenStudy (anonymous):

identical question answered here http://openstudy.com/study#/updates/55ff6651e4b0b395cadc64aa

Nnesha (nnesha):

like i said \[\sqrt{350} = 18.7\] you will get decimal answer so 350 isn't a perfect square what are the factors of 700 ?

OpenStudy (some.random.cool.kid):

please walk me through this too i dont want a straight answer

OpenStudy (some.random.cool.kid):

what factors do you mean? like single digit numbers?

Nnesha (nnesha):

yes like factors of 9 are 1 and 3 ,9 in other words 9 is divisible by 1 and 3 and 9

OpenStudy (some.random.cool.kid):

wait here are what i know

OpenStudy (some.random.cool.kid):

1, 2, 4, 5, 7, 10, 14, 20, 25, 28, 35, 50, 70? do any those work?

OpenStudy (some.random.cool.kid):

if not im not sure what other factors their are for 700

Nnesha (nnesha):

what about 100 ? we need GREATEST factor

OpenStudy (some.random.cool.kid):

oh

Nnesha (nnesha):

7 times 100 = 700 and 100 is perfect square root ?

OpenStudy (some.random.cool.kid):

yeperz 10 x 10

OpenStudy (some.random.cool.kid):

so what do I do with this -700 then now that we figured out the two multiples?

Nnesha (nnesha):

yes right so we can writeh -700 as -1*7*100 100 bec we need greatest factor and sqrt{-1} bec we can change that to i \[\huge\rm \sqrt{-1} =i\]

Nnesha (nnesha):

\[\huge\rm \frac{ -10 \pm \sqrt{-1 \times 100 \times 7} }{ 2 }\] sqrt{100} = 10 and replace sqrt{-1} with i

OpenStudy (some.random.cool.kid):

ah ha ok ok quick questions Im not sure if what I learned is the best method for a short cut but can I divide both the -10 and that sqrt by 2? and then I believe there is some final step afterwards.

Nnesha (nnesha):

first deal with the square root part then you can take out the common factor at the numerator after that you will be able to divide by 2

OpenStudy (some.random.cool.kid):

alright so whats next then?

Nnesha (nnesha):

alright so sqrt{-700}= what ??(not the decimal answer)

OpenStudy (some.random.cool.kid):

xD thats funny because I was about to put the decimal again thanks for clarifying lol. So based off of what I see you've got the 7 and 100 as your two multiples correct?

Nnesha (nnesha):

\[\huge\rm \frac{ -10 \pm \sqrt{\color{Red}{-1 \times 100 \times 7}} }{ 2 }\] that's what we got whenever you see negative sign under the square root you should know that u have to take out the negative -1 so we can convert it to i

Nnesha (nnesha):

\[\huge\rm \frac{ -10 \pm \sqrt{\color{Red}{-1 \times 100 \times 7}} }{ 2 }\] can be written as \[\frac{ -10 \pm \sqrt{-1}\sqrt{100}\sqrt{7} }{ 2}\]

OpenStudy (some.random.cool.kid):

yep i have seen that before.

Nnesha (nnesha):

now simplify sqrt{100} =? sqrt{-1}=?

OpenStudy (some.random.cool.kid):

sqrt of 100 10x10?

Nnesha (nnesha):

you can write 100 as 10 times 10 (and when we multiply same bases we should `ADD` their exponents ) so 10 times 10 = 10^2 sqrt{10^2} =?

OpenStudy (some.random.cool.kid):

wait isn't the sqrt of 10^2 the same as saying just 10?

Nnesha (nnesha):

yes right !

Nnesha (nnesha):

and sqrt{-1} =what ?

OpenStudy (some.random.cool.kid):

{-1} {-1}?

OpenStudy (some.random.cool.kid):

you write it twice right ?

Nnesha (nnesha):

hmm reread my comments

OpenStudy (some.random.cool.kid):

your adding what the sqrt{-1}?

OpenStudy (some.random.cool.kid):

sqrt -1 all i see that is would be { I }

Nnesha (nnesha):

i'm asking sqrt{-1} equal to what scroll up you will see huge latex!

Nnesha (nnesha):

I as 1 or i lol ?

OpenStudy (some.random.cool.kid):

i

OpenStudy (some.random.cool.kid):

in latex I see -1 x 100 x 7?

Nnesha (nnesha):

\[\frac{ -10 \pm 10i \sqrt{7} }{ 2 }\] now you can take out the common factor at the numerator

Nnesha (nnesha):

http://prntscr.com/8iohs8

OpenStudy (some.random.cool.kid):

yeah i said it was i

Nnesha (nnesha):

remember what is common in both terms ?